OpenStudy (anonymous):

please show me the working out of: 3(m+1) + 2(m+2) = 52

OpenStudy (anonymous):

m=9

OpenStudy (anonymous):

on both sides?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

3m+3+2m+4=52 5m=52-7 5m=45 m=9

OpenStudy (anonymous):

oh, i get it now! But were did you get the 7 from when you put 52-7 ?